OpenStudy (anonymous):

Please, step by step guide to: d/dx sinx/(1+ cosx)

OpenStudy (anonymous):

Are you there? I need you there to guide step by step to make sure that you are with me

OpenStudy (anonymous):

you have formula for this problem. that is [(sinx)' (1+ cosx) - (1+cosx)' sinx]/ (1 + cosx)^2 Now, i separate into 2 parts: numerator and denominator. numerator first: (sinx)' = cosx; (1 + cos x)' = -sinx. so, the numerator = cosx (1 + cosx) - (-sinx)(sinx)= cosx + (cosx)^2 + (sinx)^2 and we have (cosx)^2 + (sinx)^2 =1. so, the numerator = cosx +1 = 1 + cosx. combine with denominator, we have the solution is 1/ 1+ cosx. hope this help